Suppose that on the first part a 225-mile trip you averaged 55 mph. On the last part of the trip, due to traffic congestion, you only average 48 mph. If the total trip took 4 hours and 15 minutes, find the amount of time you spent traveling at each speed.
Answer by mananth(16949)

let distance of first leg be x miles
distance of second leg = 225-x miles
..
Time for first leg = x/55
time for second leg = x/48
..
Total time = 4 h 15 m = 4 1/4 hours = 17/4 hours
...
x/55 + x/48 = 17/4
48x+55x / 48*55 = 17/4
103x = 17*48*55 /4
x= (17*48*55)/(4*103)
x= 109 miles
...
109/55 = time on first leg = 1.98 hours = 2hours
Balance time on second leg = 2 h 15 m